# Classification of thin Jordan schemes

### v1

@@ Abstract.
Jordan schemes generalize association schemes in a similar way as Jordan algebras generalize the associative ones. It is well-known that association schemes of maximal rank are in one-to-one correspondence with groups (so-called thin schemes). In this paper, we classify Jordan schemes of maximal rank-to-order ratio and show that regular Jordan schemes correspond to a special class of Moufang loops, known as Ring Alternative loops.
@@ 1. Introduction
In 1959, Shah proposed [14] to use the Jordan matrix product as an algebraic tool for the analysis of block designs. This was a far-reaching generalization of the concept of an association scheme introduced in the same year by Bose and Mesner [3]. This concept was further developed and led to a flourishing research area called ”Algebraic Combinatorics” by Bannai and Ito in their famous monograph [2]. On the other hand, Shah’s idea was never noticed until Bailey published her book  [1] where she analysed the ”first” properties of the objects arising from Shah’s idea. In particular, she showed that the symmetrization of an association scheme always produces objects introduced by Shah. In 2001 Cameron called these objects Jordan schemes and posed the question whether any symmetric Jordan scheme is a symmetrization of an association scheme. The anwser was obtained in 2019 by Klin, Muzychuk and Reichard [10] who gave a negative answer to Cameron’s question and started a systematic investigation of Jordan schemes and, more generally, coherent Jordan configurations.
It is well known that all finite groups give rise to a special subclass of association schemes: the thin schemes [15]. They may be characterized as those that have a maximal rank-to-order ratio among all association schemes. In this paper, we classify all Jordan schemes that reach the maximum rank-to-order ratio. We also show that regular Jordan schemes with this property are in a one-to-one correspondence with a special class of Moufang loops.
The paper is organized as follows. Section 2 contains all necessary definitions and basic facts about coherent Jordan configurations and Jordan schemes. The third section is the central one of the paper. It is shown that the rank-to-order ratio of a JS is bounded by $3/2$ where equality happens if and only if the scheme is non-regular and thin. All such Jordan schemes are classified in Subsection 3.1. Subsection 3.2 deals with the thin regular Jordan schemes. It is proven there that there is a one-to-one correspondence between Jordan schemes of this type and Ring Alternative Moufang loops (RA-loops). In the fourth section it is shown that the Jordan schemes corresponding to RA-loops are autonomous, that is to say, they are not algebraic fusions of coherent configurations.
@@ 2.1. Binary relations and matrices
Let $\Omega$ be a non-empty finite set. A binary relation on $\Omega$ is an arbitrary subset $s\subseteq\Omega\times\Omega$. We denote by $s{\omega}:=\{\beta\in\Omega\mid(\omega,\beta)\in s\}$ an $s$-neighborhood of $\omega$. A relation is called regular if $|s\omega|$ does not depend on $\omega$. In this case, the constant $|s\omega|$ is called valency of $s$ and will be denoted as $n_{s}$. If $s$ is a regular relation of valency one, then $s{\omega}$ contains a unique element which will be denoted as $s(\omega)$, i.e., $s{\omega}=\{s(\omega)\}$. Thus $s$ is a function on $\Omega$. All functions in this paper, including permutations of ${\sf Sym}(\Omega)$, will be considered as binary relations.
Every relation $s$ admits a transposed one $s^{t}=\{(\alpha,\beta)\in\Omega^{2}\mid(\beta,\alpha)\in s\}$. Following [5] we call a relation $s\subseteq\Omega\times\Omega$ thin if $|s\omega|\leq 1\geq|s^{t}\omega|$ holds for each $\omega\in\Omega$.
We use $ab$ for the relational product of $a,b\subseteq\Omega^{2}$ and $a\star b=ab\cup ba$ for their Jordan product. The diagonal relation $1_{\Delta},\Delta\subseteq\Omega$ is defined as $1_{\Delta}:=\{(\delta,\delta)\mid\delta\in\Delta\}$. The adjacency matrix of $s\subseteq\Omega^{2}$ is the square matrix, notation $\underline{s}$, whose rows and columns are labeled by elements of $\Omega$. Its $(\alpha,\beta)$-entry $\underline{s}_{\alpha,\beta}$ is one if $(\alpha,\beta)\in s$ and zero, otherwise.
The algebra of all $\Omega\times\Omega$ matrices over a field $\mathbb{F}$ will be denoted as $M_{\Omega}(\mathbb{F})$. It is always assumed that ${\mathsf{c}har}({\mathbb{F}})\neq 2$. The standard product of two matrices $A,B\in M_{\Omega}(\mathbb{F})$ will be written as $A\cdot B$ or $AB$. Their Jordan product $\frac{1}{2}(AB+BA)$ will denoted as $A\star B$. The Schur-Hadamard (entry-wise) product will be denoted as $A\circ B$. The identity matrix and all-one matrix will be denoted as $I_{\Omega}$ and $J_{\Omega}$, respectively. The linear span of any set of matrices ${\mathcal{S}}\subseteq M_{\Omega}({\mathbb{F}})$ will be denoted as $\langle{\mathcal{S}}\rangle$.
@@ 2.2. Rainbow algebras
Consider a vector subspace ${\mathcal{A}}\subseteq M_{\Omega}(\mathbb{F})$ which is closed with respect to t and $\circ$ and contains $I_{\Omega},J_{\Omega}$. Every such subspace ${\mathcal{A}}\subseteq M_{\Omega}(\mathbb{F})$ has a uniquely determined standard basis — the basis consisting of the minimal $\circ$-idempotents of the algebra $({\mathcal{A}},\circ)$. These are $0,1$-matrices $A_{1},...,A_{r}\in{\mathcal{A}}$ that are pairwise $\circ$-orthogonal. Each $A_{i}$ is the adjacency matrix of a unique binary relation $s_{i}\subseteq\Omega^{2}$, i.e., $A_{i}=\underline{s}_{i},i=1,...,r$. The set $S=\{s_{1},...,s_{r}\}$ forms a partition of $\Omega^{2}$. This partition satisfies the following conditions:
@@
Following [8] we call a pair $\mathfrak{S}=(\Omega,S)$ a rainbow if it satisfies the above conditions. The elements of $S$ are called basic relations of the rainbow. The numbers $|\Omega|$ and $|S|$ are called the order and the rank of the rainbow; it has rank one if and only if $|\Omega|=1$. If the point set $\Omega$ has at least two elements, then $|S|\geq 2$ where the equality holds if and only if $S=\{1_{\Omega},\Omega^{2}\setminus 1_{\Omega}\}$. These rainbows are called trivial. The adjacency matrix of a rainbow ${\mathfrak{S}}=(\Omega,S)$ is the $\Omega\times\Omega$ matrix $A({\mathfrak{S}})$ defined via $A({\mathfrak{S}})_{\alpha,\beta}=S(\alpha,\beta)$ where $S(\alpha,\beta)$ is a unique basic relation containing $(\alpha,\beta)\in\Omega^{2}$.
The rainbow of maximal rank $|\Omega|^{2}$ is called discrete. Each basic relation of a discrete rainbow contains a unique ordered pair.
In what follows, we say that a subspace ${\mathcal{A}}\subseteq M_{\Omega}(\mathbb{F})$ is a rainbow algebra if it contains $I_{\Omega},J_{\Omega}$ and is closed with respect to $^{t}$$\circ$. We say that $\mathcal{A}$} is symmetric if every $A\in{\mathcal{A}}$ symmetric; homogeneous if $\langle I_{\Omega}\rangle\circ{\mathcal{A}}\subseteq\langle I_{\Omega}\rangle$; regular if $\langle J_{\Omega}\rangle\cdot{\mathcal{A}}\subseteq\langle J_{\Omega}\rangle$.
Given a rainbow $\mathfrak{S}=(\Omega,S)$, the linear span $\mathbb{F}[S]:=\langle\underline{s}\rangle_{s\in S}\subseteq M_{\Omega}(\mathbb{F})$ is a rainbow algebra. Thus, there exists a one-to-one correspondence between rainbows and rainbow subalgebras of $M_{\Omega}(\mathbb{F})$. The statement below collects some properties of rainbow algberas.
@@
Let $\mathfrak{S}=(\Omega,S)$ be a rainbow and ${\mathcal{A}}=\langle\underline{s}\rangle_{s\in S}$ the corresponding rainbow algebra. Then
@@
In what follows, we call a rainbow symmetric, homogeneous, or regular depending on which of the properties mentioned above it satisfies. Notice that the regularity of a rainbow always implies its homogeneity but not vice versa. A rainbow will be called thin if all its basic relations are thin.
Given two rainbows $\mathfrak{S}=(\Omega,S)$ and $\mathfrak{S^{\prime}}=(\Omega,S^{\prime})$, we say that $\mathfrak{S}$ is a fusion of $\mathfrak{S^{\prime}}$ (equivalently, $\mathfrak{S^{\prime}}$ is a fission of $\mathfrak{S}$) if $\mathbb{F}[S]\subseteq\mathbb{F}[S^{\prime}]$.
Two rainbows $\mathfrak{S}=(\Omega,S)$ and $\mathfrak{S^{\prime}}=(\Omega^{\prime},S^{\prime})$ are called (combinatorially) isomorphic if there exists a bijection $f:\Omega\rightarrow\Omega^{\prime}$ satisfying $f(S):=\{f(s)\mid s\in S\}=S^{\prime}$ where $f(s):=\{(f(\alpha),f(\beta))\mid(\alpha,\beta)\in s\}$.
@@ 2.3. Coherent Jordan algebras
A rainbow algebra is called coherent algebra (see [8]) if it is closed with respect to matrix multiplication $\cdot$. Similarly, a rainbow algebra will be called a coherent Jordan algebra (coherent J-algebra, for short) if it is closed with respect to Jordan multiplication $\star$.
Let $\mathfrak{S}=(\Omega,S)$ be a rainbow, and let ${\mathcal{A}}:=\langle\underline{s}\rangle_{s\in S}$ be the corresponding rainbow algebra. We say that $\mathfrak{S}$ is a coherent configuration (respectively, a J-configuration) if $\mathcal{A}$} is a coherent algebra (respectively, a J-algebra). In both cases, the algebra $\mathcal{A}$} will be referred to as adjacency algebra of $\mathfrak{S}$. In what follows, we are going to use the abbreviations CC and JC for coherent configurations and coherent J-configurations, respectively.
If $\mathfrak{S}=(\Omega,S)$ is a coherent configuration, then for any pair $s,r\in S$ we have $\underline{s}\cdot\underline{r}=\sum_{t}p_{s,r}^{t}\underline{t}$ for suitable nonnegative integers $p_{s,r}^{t}$. These numbers are called structures constants of $\mathfrak{S}$. Due to their combinatorial meaning $p_{s,r}^{t}=|\alpha s\cap\beta r^{t}|,(\alpha,\beta)\in t$, they are also called the intersection numbers of $\mathfrak{S}$.
Similarly, one can define the intersection numbers for coherent J-configuration: $p_{\{s,r\}}^{t}=|\alpha s\cap\beta r^{t}|+|\alpha r\cap\beta s^{t}|,(\alpha,\beta)\in t$. Notice that in this case we write them as $p_{\{s,r\}}^{t}$ stressing the fact that $\star$ is commutative. They also appear as the structure constants for the Jordan product $\underline{s}\star\underline{r}=\frac{1}{2}\sum_{t}p_{\{s,r\}}^{t}\underline{t}$.
Recall that a homogeneous coherent configuration is called an association scheme (see [15]). Following this pattern, we will call a homogeneous coherent J-configuration a Jordan scheme. While it is well known that association schemes are always regular, Jordan schemes do not possess this property in general. It was shown in [12] that a symmetric JS is always regular. Regarding non-symmetric Jordan schemes, they are not necessarily regular, but have a very special structure — see [12, Proposition 2.2]. Examples of non-regular Jordan schemes also appear in Subsection 3.1 of this paper.
@@ 2.4. Coherent J-closure
The coherent J-closure generalizes the well-known concept of the coherent closure of $\mathcal{S}$ which is the intersection of all coherent algebras containing $\mathcal{S}$. It is often called the WL-closure, since Weisfeiler and Leman were the authors of the algorithm which computes it efficiently. For this reason, we are going to denote it as $\operatorname{WL}(\mathcal{S})$. Notice that we always have $J(\mathcal{S})\subseteq\operatorname{WL}(\mathcal{S})$.

### v2

@@ Abstract.
Jordan schemes generalize association schemes in a manner similar to how Jordan algebras generalize associative algebras. It is well known that association schemes of maximum rank are in one-to-one correspondence with groups (so-called thin schemes). In this paper, we classify Jordan schemes of maximum rank-to-order ratio and show that regular Jordan schemes of such type correspond to a special class of Moufang loops known as ring-alternative loops.
Key words and phrases:
Jordan algebra, thin Jordan scheme, coherent configuration, Moufang loop, RA-loop, autonomous
1991 Mathematics Subject Classification:
17C50, 20N05, 05E30, 05E16, 16W10
@@ 1. Introduction
In 1959, Shah proposed [16] to use the Jordan matrix product as an algebraic tool for the analysis of block designs. This was a far-reaching generalization of the concept of an association scheme introduced in the same year by Bose and Mesner [3]. The concept was further developed and led to a flourishing research area called “Algebraic Combinatorics” by Bannai and Ito in their famous monograph [2]. Nevertheless, Shah’s idea went virtually unnoticed until Bailey published her book  [1] where she analyzed the “first” properties of objects arising from Shah’s idea. In particular, she showed that the symmetrization of an association scheme always produces objects introduced by Shah. In 2001, Cameron called these objects Jordan schemes and posed the question, “Is every symmetric Jordan scheme a symmetrization of an association scheme?” In 2019, Klin, Muzychuk and Reichard gave a negative answer to Cameron’s question and started a systematic investigation of Jordan schemes and, more generally, coherent Jordan configurations [11].
It is well known that all finite groups form a special subclass of association schemes, called thin schemes [19]. Such schemes may be characterized as those that have a maximum rank-to-order ratio among all association schemes. In this paper, we classify all Jordan schemes that attain maximum rank-to-order ratio. We further show that regular Jordan schemes with this property are in a one-to-one correspondence with a special class of Moufang loops.
The paper is organized as follows. Section 2 contains all necessary definitions and basic facts about coherent Jordan configurations and Jordan schemes. Section 3 is the most central part of our paper. There we show that the rank-to-order ratio of a Jordan scheme is bounded by $3/2$ with equality occurring if and only if the scheme is thin and nonregular. All such Jordan schemes are classified in Subsection 3.1. Subsection 3.2 treats thin regular Jordan schemes. There we establish a one-to-one correspondence between such schemes and ring-alternative Moufang loops (RA-loops). Finally, in Section 4 we show that Jordan schemes corresponding to RA-loops are autonomous, that is to say, they are not algebraic fusions of coherent configurations. Thus we describe the first infinite family of autonomous Jordan schemes. Previously, the only known autonomous Jordan scheme was the one on $15$ points described in [14] (see also [11, Section 4]) related to one of the two Siamese $\mathrm{STS}(15)$ constructed in [13, Section 4]. The fact that this Jordan scheme is autonomous was established using the COCO2P package [12] in GAP [6] and Ziv-Av’s catalog of small coherent configurations [20].
@@ 2.1. Binary relations and matrices
Let $\Omega$ be a nonempty finite set. A binary relation on $\Omega$ is an arbitrary subset $s\subseteq\Omega\times\Omega$. We refer to the set $s{\omega}:=\{\beta\in\Omega\mid(\omega,\beta)\in s\}$ as the $s$-neighborhood of $\omega\in\Omega$. A relation is called regular if $|s\omega|$ does not depend on $\omega$. In such case, we refer to the constant $n_{s}:=|s\omega|$ as the valency of $s$. If $s$ is a regular relation of valency one, then $s{\omega}$ contains a unique element which we denote as $s(\omega)$, i.e. $s{\omega}=\{s(\omega)\}$. Thus $s$ is a function on $\Omega$. All functions in this paper, including permutations from ${\sf Sym}(\Omega)$, will be considered as binary relations.
Every relation $s$ admits a transposed relation $s^{t}=\{(\alpha,\beta)\in\Omega^{2}\mid(\beta,\alpha)\in s\}$. Throughout, we will denote the unary operator that transposes relations by $\tau$, that is, $\tau:s\mapsto s^{t}$. Following [5] we call a relation $s\subseteq\Omega\times\Omega$ thin if $|s\omega|\leq 1\geq|s^{t}\omega|$ holds for each $\omega\in\Omega$.
We write $ab$ for the relational product of $a,b\subseteq\Omega^{2}$ and $a\star b=ab\cup ba$ for their Jordan product. For $\Delta\subseteq\Omega$, the diagonal relation $1_{\Delta}$ is defined by $1_{\Delta}:=\{(\delta,\delta)\mid\delta\in\Delta\}$. The adjacency matrix of $s\subseteq\Omega^{2}$ is the square matrix, denoted $\underline{s}$, whose rows and columns are labeled by the elements of $\Omega$ and whose $(\alpha,\beta)$-entry $\underline{s}_{\alpha,\beta}$ equals one if $(\alpha,\beta)\in s$ and zero otherwise.
The algebra of all $\Omega\times\Omega$ matrices over a field $\mathbb{F}$ will be denoted as $M_{\Omega}(\mathbb{F})$. It is always assumed that ${\mathsf{c}har}({\mathbb{F}})\neq 2$. The standard product of two matrices $A,B\in M_{\Omega}(\mathbb{F})$ will be written as $A\cdot B$ or $AB$. while their Jordan product $\frac{1}{2}(AB+BA)$ will be denoted by $A\star B$. The Schur-Hadamard (or entry-wise) product will be denoted as $A\circ B$. The identity matrix and all-ones matrix will be denoted as $I_{\Omega}$ and $J_{\Omega}$, respectively. The linear span of any set of matrices ${\mathcal{S}}\subseteq M_{\Omega}({\mathbb{F}})$ will be denoted by $\mathbb{F}[\mathcal{S}]$. In the case where $\mathcal{S}$ is given as a family $\{A_{i}\}_{i\in I}$, we also write $\langle A_{i}\rangle_{i\in I}$ for $\mathbb{F}[\mathcal{S}]$.
@@ 2.2. Rainbow algebras
Consider a vector subspace ${\mathcal{A}}\subseteq M_{\Omega}(\mathbb{F})$ which is closed with respect to $\tau$ and $\circ$ and contains $I_{\Omega},J_{\Omega}$. Every such subspace ${\mathcal{A}}\subseteq M_{\Omega}(\mathbb{F})$ has a uniquely determined standard basis, namely the basis consisting of all minimal $\circ$-idempotents of the algebra $({\mathcal{A}},\circ)$. These are $\{0,1\}$-matrices $A_{1},\dots,A_{r}\in{\mathcal{A}}$ that are pairwise $\circ$-orthogonal. Each $A_{i}$ is the adjacency matrix of a unique binary relation $s_{i}\subseteq\Omega^{2}$, i.e. $A_{i}=\underline{s}_{i},i=1,\dots,r$. The set $S=\{s_{1},\dots,s_{r}\}$ forms a partition of $\Omega^{2}$ that satisfies the following two conditions:
@@
Following [8] we call a pair $\mathfrak{S}=(\Omega,S)$ a rainbow if it satisfies these conditions. The elements of $S$ are called basic relations of the rainbow. The cardinalities $|\Omega|$ and $|S|$ are called the order and rank of the rainbow, respectively. Observe that a rainbow has rank one if and only if $|\Omega|=1$. For $|\Omega|\geq 2$ one has $|S|\geq 2$, where equality is attained if and only if $S=\{1_{\Omega},\,\Omega^{2}\setminus 1_{\Omega}\}$. Such rainbows are called trivial. The adjacency matrix of a rainbow ${\mathfrak{S}}=(\Omega,S)$ is the $\Omega\times\Omega$ matrix $A({\mathfrak{S}})$ defined via $A({\mathfrak{S}})_{\alpha,\beta}=S(\alpha,\beta)$ where $S(\alpha,\beta)$ is the unique basic relation containing $(\alpha,\beta)\in\Omega^{2}$.
A rainbow of maximum rank $|\Omega|^{2}$ is called discrete. Each basic relation of a discrete rainbow contains a unique ordered pair.
In what follows, we call a subspace ${\mathcal{A}}\subseteq M_{\Omega}(\mathbb{F})$ a rainbow algebra if it contains $I_{\Omega},J_{\Omega}$ and is closed with respect to $\tau$ and $\circ$. We say $\mathcal{A}$} is symmetric if every $A\in{\mathcal{A}}$ symmetric, i.e. $A^{t}=A$ for every $A\in{\mathcal{A}}$. We call $\mathcal{A}$} homogeneous if $\langle I_{\Omega}\rangle\circ{\mathcal{A}}\subseteq\langle I_{\Omega}\rangle$ and regular if $\langle J_{\Omega}\rangle\cdot{\mathcal{A}}\subseteq\langle J_{\Omega}\rangle$.
Given a rainbow $\mathfrak{S}=(\Omega,S)$, the linear span $\mathbb{F}[S]:=\langle\underline{s}\rangle_{s\in S}\subseteq M_{\Omega}(\mathbb{F})$ is a rainbow algebra. Thus, there exists a one-to-one correspondence between rainbows and rainbow subalgebras of $M_{\Omega}(\mathbb{F})$. The statement below collects some properties of rainbow algebras.
@@
Let $\mathfrak{S}=(\Omega,S)$ be a rainbow, and let ${\mathcal{A}}=\langle\underline{s}\rangle_{s\in S}$ be the corresponding rainbow algebra. Then
@@
In what follows, we call a rainbow symmetric, homogeneous, or regular depending on which of those properties its corresponding algebra satisfies. Note that regularity of a rainbow always implies homogeneity but not vice versa. A rainbow will be called thin if all of its basic relations are thin.
Given two rainbows $\mathfrak{S}=(\Omega,S)$ and $\mathfrak{S^{\prime}}=(\Omega,S^{\prime})$, we say that $\mathfrak{S}$ is a fusion of $\mathfrak{S^{\prime}}$ (equivalently, $\mathfrak{S^{\prime}}$ is a fission of $\mathfrak{S}$) if $\mathbb{F}[S]\subseteq\mathbb{F}[S^{\prime}]$. In this case, we also write $S\sqsubseteq S^{\prime}$.
Two rainbows $\mathfrak{S}=(\Omega,S)$ and $\mathfrak{S^{\prime}}=(\Omega^{\prime},S^{\prime})$ are said to be (combinatorially) isomorphic if there exists a bijection $f:\Omega\rightarrow\Omega^{\prime}$ satisfying $f(S)=S^{\prime}$, where $f(S):=\{f(s)\mid s\in S\}$ and $f(s):=\{(f(\alpha),f(\beta))\mid(\alpha,\beta)\in s\}$.
@@ 2.3. Coherent Jordan algebras
A rainbow algebra is called a coherent algebra (see [8]) if it is closed with respect to standard matrix multiplication. Similarly, a rainbow algebra will be called a coherent Jordan algebra (coherent J-algebra, for short) if it is closed with respect to Jordan multiplication $\star$.
Let $\mathfrak{S}=(\Omega,S)$ be a rainbow, and let ${\mathcal{A}}:=\langle\underline{s}\rangle_{s\in S}$ be its corresponding rainbow algebra. We say $\mathfrak{S}$ is a coherent configuration (resp. J-configuration) if $\mathcal{A}$} is a coherent algebra (resp. J-algebra). In both cases, the algebra $\mathcal{A}$} will be referred to as the adjacency algebra of $\mathfrak{S}$.
If $\mathfrak{S}=(\Omega,S)$ is a coherent configuration, then for any pair $s,r\in S$ we have $\underline{s}\cdot\underline{r}=\sum_{t}p_{s,r}^{t}\underline{t}$ for suitable nonnegative integers $p_{s,r}^{t}$. We refer to these $p_{s,r}^{t}$ as structure constants of $\mathfrak{S}$. Due to their combinatorial meaning $p_{s,r}^{t}=|\alpha s\cap\beta r^{t}|,(\alpha,\beta)\in t$, they are also called intersection numbers of $\mathfrak{S}$.
Similarly, one can define intersection numbers $p_{\{s,r\}}^{t}$ for a coherent J-configuration as follows: $p_{\{s,r\}}^{t}=|\alpha s\cap\beta r^{t}|+|\alpha r\cap\beta s^{t}|,(\alpha,\beta)\in t$. These also appear as structure constants for the Jordan product $\underline{s}\star\underline{r}=\frac{1}{2}\sum_{t}p_{\{s,r\}}^{t}\underline{t}$. (Note that our notation $p_{\{s,r\}}^{t}$ reflects the fact that $\star$ is commutative.)
Recall that a homogeneous coherent configuration is called an association scheme (see [19]). Following this pattern, we will call a homogeneous coherent J-configuration a Jordan scheme. While it is well known that association schemes are always regular, Jordan schemes do not possess this property in general. It was shown in [14] that a symmetric Jordan scheme is always regular. Regarding nonsymmetric Jordan schemes, they are not necessarily regular, but have a very special structure — see [14, Proposition 2.2]. Examples of nonregular Jordan schemes also appear in Subsection 3.1 of this paper.
@@ 2.4. Coherent J-closure
The coherent J-closure generalizes the well-known concept of the coherent closure of $\mathcal{S}$ which is the intersection of all coherent algebras containing $\mathcal{S}$. It is often called the WL-closure in honor of Weisfeiler and Leman [17], who designed an algorithm for its computation. For this reason, we will denote this closure by $\operatorname{WL}(\mathcal{S})$. Observe that we always have $J(\mathcal{S})\subseteq\operatorname{WL}(\mathcal{S})$. If $S$ is a set of binary relations on $\Omega$, then we write $J(S)$ and $\operatorname{WL}(S)$ for the underlying rainbows of $J(\{\underline{s}\}_{s\in\mathcal{S}})$ and $\operatorname{WL}(\{\underline{s}\}_{s\in\mathcal{S}})$, respectively.